# Token-Importance Guided Direct Preference Optimization

### v2

@@ 1 Introduction
Large Language Models (LLMs) have shown proficiency in Natural Language Processing (NLP) (Gao et al., 2025), logical reasoning (Xie et al., 2025a), and code generation (Xu et al., 2025), emerging as a focal point of recent research. However, as models may generate outputs inconsistent with intended purposes or ethical standards, human preference alignment aims to ensure that LLMs adhere to human values (Liu et al., 2023), producing beneficial and harmless content. Against this backdrop, Reinforcement Learning from Human Feedback (RLHF) has become a prevailing approach for achieving alignment (Hong et al., 2024; Hu et al., 2025). It leverages human-annotated preference data to train reward models and fine-tunes LLMs using Reinforcement Learning (RL) methods (Wang et al., 2023b) like Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@
Motivated by these challenges, researchers have proposed token-level variants of DPO, aiming to decompose preference alignment into fine-grained contributions ((Zeng et al., 2024; Xie et al., 2025a; Zhong et al., 2024)). However, achieving true fine-grained alignment requires addressing a core challenge: We not only need to accurately identify the key tokens that have a decisive impact on human preferences, but also need a subtle optimization objective to guide the model to adjust its preference (Li et al., 2025). Nevertheless, existing token-level methods fall short in dealing with this challenge for two reasons. First, their approaches to identifying key tokens often rely on biased probability proxies (Liu et al., 2024a) or overly simplified weighting schemes (Lin et al., 2024). Second, the optimization, they still inherit the binary comparison framework of DPO, simply distinguishing between “good” and “bad” samples (Meng et al., 2024). Such coarse-grained supervision signals cannot finely shape the model’s generation behavior in a continuous semantic space.
@@ 2 Related Work
From Sequence-Level to Token-Level Achieving fine-grained alignment requires the model not only to distinguish the quality of the entire sequence, but also to understand and precisely control the key morphemes that constitute the semantics of the sequence. Existing sequence-level techniques often lead to a decrease in generation diversity because they ignore the importance differences among tokens (Feng et al., 2025). These limitations have spurred researchers’ research into step (Xie et al., 2024) or token-level (Rafailov et al., 2024) alignment mechanisms, seeking to address the granularity mismatch between coarse-grained sequence rewards and fine-grained token contributions (Xi et al., 2024). To address the significant decline in model generation diversity, TDPO (Zeng et al., 2024) reanalyzed and optimized the entire alignment process from the token-level perspective. An additional limitation of RLHF and DPO lies in the fact that rewards are only assigned to the final token, with all other tokens receiving no learning rewards (Zhong et al., 2024). Meanwhile, Xie et al. (2025a) proposed a correlation between the frequency of specific tokens and model performance, which inspires us to consider reassigning token weights. In a related vein, Liu et al. (2024a) estimates token importance weights using prediction probability differences. Nevertheless, this probabilistic weighting scheme is prone to bias when contrastive models produce inconsistent outputs or fail to capture subtle semantic nuances of human preferences. In contrast, our approach employs a hybrid strategy that combines causal gradient attribution with a stabilizing Gaussian prior to estimate importance (Ballout et al., 2024). By focusing on actual gradient impacts, our method enhances alignment precision over probabilistic proxies, while the prior distribution ensures robustness against noisy gradient signals.
@@ 4.2 Hybrid Weighting Mechanism for Token Importance
Finally, to ensure training stability and robustness against noise in gradient estimates, we post-process these raw scores to derive the final weights $w_{t}$. As implemented in our code, this involves a mixed strategy. First, the raw scores $\mathcal{I}=\{I_{1},\dots,I_{T-1}\}$ are normalized by their sum to form a distribution $\mathcal{I}_{\text{norm}}$. Second, we define a Gaussian-shaped prior distribution $\mathcal{P}_{\text{prior}}$ centered on the sequence, which assigns higher baseline importance to tokens in the middle. (Detailed settings for $\mathcal{P}_{\text{prior}}$ can be found in Appendix B.6.) The final weight vector $W$ is a convex combination of these two distributions, controlled by a hyperparameter $\lambda\in[0,1]$:
@@ 4.4 TI-DPO Objective and Theoretical Analysis
Lemma 1 (Variance Reduction).
$\sigma^{2}_{\text{TI-DPO}}<\sigma^{2}_{\text{DPO}}$ (14)
Theorem 2 (Tighter Loss Bound).
Theorem 3 (Superiority of Optimal Policy).
$\mathbb{E}_{y\sim\pi_{TI-DPO}}[r^{*}(x,y)]\geq\mathbb{E}_{y\sim\pi_{DPO}}[r^{*}(x,y)]+\delta,$ (16)
5 Experiments
5.1 Experimental Settings
Dataset and base settings: The benchmarks we use include knowledge-based tasks (MMLU (Hendrycks et al., 2020)), mathematical reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021)), instruction-following (IFEval (Zhou et al., 2023)), and code generation (HumanEval (Chen et al., 2021)). Additionally, TruthfulQA (Lin et al., 2021) detects the authenticity of the model’s answers through adversarial questions.
Comparative algorithm: We compared the TI-DPO with baseline alignment methods such as SFT, DPO, IPO(Azar et al., 2024), KTO (Ethayarajh et al., 2024), SimPO (Meng et al., 2024), TDPO (Zeng et al., 2024), CPO (Feng et al., 2025), TPO (Saeidi et al., 2024), TIS-DPO (Liu et al., 2024a), Logic-RL (Xie et al., 2025b), cDPO (Lin et al., 2024) and GRPO (Shao et al., 2024). We select three models (Llama-3.2-3B (Grattafiori et al., 2024), Llama-3.1-8B (Grattafiori et al., 2024), Mistral-7B-v0.3 (Jiang et al., 2023)) as baselines.
@@ 5.1 Experimental Settings
5.2 Performance Comparison
(a) TruthfulQA Performance vs. Training Steps
(b) IFEval Performance vs. Training Steps
@@ 5.4 Additional Experiments
Case Study: A case study on a medical query (see Appendix C.1 for details) demonstrates that, given the user prompt “I have a headache, what should I do?”, TI-DPO effectively assigns higher importance to safety-critical tokens (e.g., “medical attention”, “promptly”) in preferred responses, while penalizing risky suggestions (e.g., “painkillers”, “casually”) in non-preferred ones. Additionally, there are two other cases in Appendix C.2.
Pearson correlation coefficient: To investigate the effectiveness of our hybrid weighting mechanism, we also conducted Pearson correlation coefficient analysis, with full results and methodology presented in Appendix B.2.
Robustness and Generation Diversity: We validate the robustness and generative diversity of TI-DPO, which can be seen in Appendix B.4.
Sensitivity of Hyperparameters: We conducte sensitivity analyses for the weight-mixing parameter $\lambda$ and KL weight $\alpha$ in Appendix B.6. And we have provided the specific values of the hyperparameters in this project, which was also shown in Appendix B.6.
@@ 6 Conclusion
We introduce TI-DPO, an optimization framework that effectively bridges the alignment gap between LLMs and human value systems. By introducing a mixed weight calculated collaboratively by gradient attribution and Gaussian prior, TI-DPO effectively overcomes the limitations of traditional DPO methods at the token level and their sensitivity to noise. On this basis, the triplet loss structure provides more refined guidance for model optimization. Theorem 2 and Theorem 3 theoretically illustrate the superiority of TI-DPO over DPO. The effectiveness of TI-DPO is unequivocally demonstrated through extensive experimentation. Our method achieves a state-of-the-art average score of 62.3 across a diverse suite of benchmarks, outperforming all baseline methods. This superiority is particularly evident in complex alignment tasks: On HumanEval, TruthfulQA and IFEval, TI-DPO scores 67.0, 62.0 and 75.7 respectively, significantly surpassing strong contenders like GRPO and traditional DPO.
As for the limitations, despite its effectiveness in fine-grained alignment, TI-DPO entails a computational overhead during training and performs slightly below sequence-level baselines on holistic reasoning tasks. Future work will focus on integrating our token-importance mechanism with group-based optimization methods like GRPO to bridge this gap and further enhance reasoning capabilities. More statements can be found in Appendix D.
In summary, TI-DPO demonstrates outstanding performance in enhancing the output quality and reliability of LLMs through its innovative design.
@@ A.4 Gradient Analysis of Loss
Table A1: Distribution of Token Importance Weight, Performance Improvement, and Sample-level Pearson Correlation Coefficient in Each Task
@@
Appendix B Additional Experimental Results
We present some additional explanations and experimental results.
B.1 Distribution of Weights
Figure B1 visualizes the distribution of importance weights assigned to tokens in different tasks by TI-DPO with histograms and box plots, intuitively explaining how the importance mechanism dynamically focuses on key tokens according to task characteristics. The y-axis of the histogram is token frequency, which represents the number of occurrences of tokens with varying importance. The box plots in the upper right corner are used to visualize the distribution of token frequencies for different weights, where the red line represents the median, reflecting the concentration of token frequencies. In the GSM8K and GPQA datasets, with only a few key symbols crucial for answering, the importance weights mostly concentrate in the interval of $[0.2,0.5]$ approximately. For the TruthfulQA and IFEval datasets, token distribution is concentrated in the $[0.6,0.8]$ weight interval. The MMLU and HumanEval datasets cover a wide range of content, and the model assigns diverse weights to various tokens.
Figure B1: Distribution patterns of gradient-based token importance weights in six benchmark tasks (GSM8K, TruthfulQA, MMLU, GPQA, HumanEval, IFEval
B.2 Pearson Correlation Coefficient
Table A1 presents multiple metrics for six tasks, namely GSM8K, GPQA, TruthfulQA, IFEval, MMLU, and HumanEval, including Q1, Q2 (median), Q3, average weight (mid-point of the interval), performance improvement $\Delta_{\text{ACC}}$ (%), and sample-level Pearson $r$ (Top 5 weights $vs$ accuracy). Among them, the “sample-level Pearson rr” is calculated based on the average weights of the Top-5 tokens for 100 questions in each task and the correctness of the answers to these questions (binary values), which reflects the microscopic internal correlation. The performance improvement $\Delta_{\text{ACC}}$ is calculated from the average accuracy gain of TI-DPO compared to DPO. The calculation process is divided into the calculation of microscopic and macroscopic Pearson correlation coefficients:
@@ B.2 Pearson Correlation Coefficient
B.3 Experimental Results Across Base Models
@@
B.4 Robustness and Generative Diversity
@@
B.5 Convergence Based on Theorem 2
@@
B.6 Sensitivity Analysis and Hyperparameters
@@
For each token position $t\in[0,T-1]$, the unnormalized value is calculated as $\mathcal{P}_{\text{prior}}(t)=\exp\left(-\frac{1}{2}\left(\frac{t-\mu}{\sigma}\right)^{2}\right).$ Here, we specifically chose $\mu=(T-1)/2$ and $\sigma=T/4$ as a robust geometric heuristic. Since approximately 95% of the mass of a Gaussian distribution lies within $\pm 2\sigma$, setting $4\sigma\approx T$ ensures the prior effectively spans the entire sequence context without being too narrow or too flat.
@@ C.1 Analysis of Medical-related Case
Figure B2: Case demo of responses to prompt “I have a headache, what should I do?”. Left: Preferred case. Middle: Intermediate case. Right: Non-preferred case. The darker color indicates higher weight.

### v3

# arXiv:2505.19653v3[cs.AI] 02 Mar 2026

@@ 1 Introduction
Large Language Models (LLMs) have shown proficiency in Natural Language Processing (NLP) (Gao et al., 2025), logical reasoning (Xie et al., 2025), and code generation (Xu et al., 2025), emerging as a focal point of recent research. However, as models may generate outputs inconsistent with intended purposes or ethical standards, human preference alignment aims to ensure that LLMs adhere to human values (Liu et al., 2023), producing beneficial and harmless content. Against this backdrop, Reinforcement Learning from Human Feedback (RLHF) has become a prevailing approach for achieving alignment (Hong et al., 2024; Hu et al., 2025). It leverages human-annotated preference data to train reward models and fine-tunes LLMs using Reinforcement Learning (RL) methods (Wang et al., 2023b) like Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@
Motivated by these challenges, researchers have proposed token-level variants of DPO, aiming to decompose preference alignment into fine-grained contributions ((Zeng et al., 2024; Xie et al., 2025; Zhong et al., 2024)). However, achieving true fine-grained alignment requires addressing a core challenge: We not only need to accurately identify the key tokens that have a decisive impact on human preferences, but also need a subtle optimization objective to guide the model to adjust its preference (Li et al., 2025). Nevertheless, existing token-level methods fall short in dealing with this challenge for two reasons. First, their approaches to identifying key tokens often rely on biased probability proxies (Liu et al., 2024a) or overly simplified weighting schemes (Lin et al., 2024). Second, the optimization, they still inherit the binary comparison framework of DPO, simply distinguishing between “good” and “bad” samples (Meng et al., 2024). Such coarse-grained supervision signals cannot finely shape the model’s generation behavior in a continuous semantic space.
@@ 2 Related Work
From Sequence-Level to Token-Level Achieving fine-grained alignment requires the model not only to distinguish the quality of the entire sequence, but also to understand and precisely control the key morphemes that constitute the semantics of the sequence. Existing sequence-level techniques often lead to a decrease in generation diversity because they ignore the importance differences among tokens (Feng et al., 2025). These limitations have spurred researchers’ research into step (Xie et al., 2024) or token-level (Rafailov et al., 2024) alignment mechanisms, seeking to address the granularity mismatch between coarse-grained sequence rewards and fine-grained token contributions (Xi et al., 2024). To address the significant decline in model generation diversity, TDPO (Zeng et al., 2024) reanalyzed and optimized the entire alignment process from the token-level perspective. An additional limitation of RLHF and DPO lies in the fact that rewards are only assigned to the final token, with all other tokens receiving no learning rewards (Zhong et al., 2024). Meanwhile, Xie et al. (2025) proposed a correlation between the frequency of specific tokens and model performance, which inspires us to consider reassigning token weights. In a related vein, Liu et al. (2024a) estimates token importance weights using prediction probability differences. Nevertheless, this probabilistic weighting scheme is prone to bias when contrastive models produce inconsistent outputs or fail to capture subtle semantic nuances of human preferences. In contrast, our approach employs a hybrid strategy that combines causal gradient attribution with a stabilizing Gaussian prior to estimate importance (Ballout et al., 2024). By focusing on actual gradient impacts, our method enhances alignment precision over probabilistic proxies, while the prior distribution ensures robustness against noisy gradient signals.
@@ 4.2 Hybrid Weighting Mechanism for Token Importance
Finally, to ensure training stability and robustness against noise in gradient estimates, we post-process these raw scores to derive the final weights $w_{t}$. As implemented in our code, this involves a mixed strategy: (1) First, the raw scores $\mathcal{I}=\{I_{1},\dots,I_{T-1}\}$ are normalized by their sum to form a distribution $\mathcal{I}_{\text{norm}}$. (2) Then we define a Gaussian-shaped prior distribution $\mathcal{P}_{\text{prior}}$ centered on the sequence, which assigns higher baseline importance to tokens in the middle. For a sequence of length $T$ and each token position $t\in[0,T-1]$, the unnormalized value is calculated as:
$\mathcal{P}_{\text{prior}}(t)=\exp\left(-\frac{1}{2}\left(\frac{t-\mu}{\sigma}\right)^{2}\right),$ (7)
where we heuristically set the mean $\mu=(T-1)/2$ and the standard deviation $\sigma=T/4$. Since approximately 95% of the mass of a Gaussian distribution lies within $\pm 2\sigma$, setting $4\sigma\approx T$ ensures the prior effectively spans the entire sequence context without being too narrow or too flat.
The final weight vector $W$ is a convex combination of these two distributions, controlled by a hyperparameter $\lambda\in[0,1]$:
@@ 4.4 TI-DPO Objective and Theoretical Analysis
Lemma 1 (Variance Reduction).
$\sigma^{2}_{\text{TI-DPO}}<\sigma^{2}_{\text{DPO}}$ (15)
Theorem 2 (Tighter Loss Bound).
Theorem 3 (Superiority of Optimal Policy).
$\mathbb{E}_{y\sim\pi_{\text{TI-DPO}}}[r^{*}(x,y)]\geq\mathbb{E}_{y\sim\pi_{\text{DPO}}}[r^{*}(x,y)]+\delta,$ (17)
5 Experiments
5.1 Experimental Settings
Datasets and base settings. The benchmarks we use include knowledge-based tasks (MMLU (Hendrycks et al., 2020)), mathematical reasoning (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021)), instruction-following (IFEval (Zhou et al., 2023)), and code generation (HumanEval (Chen et al., 2021)). Additionally, TruthfulQA (Lin et al., 2021) detects the authenticity of the model’s answers through adversarial questions. For other detailed hyperparameter settings, please refer to Appendix B.5.
Comparative algorithm. We compared the TI-DPO with baseline alignment methods such as SFT, DPO, IPO(Azar et al., 2024), KTO (Ethayarajh et al., 2024), SimPO (Meng et al., 2024), TDPO (Zeng et al., 2024), CPO (Feng et al., 2025), TPO (Saeidi et al., 2024), TIS-DPO (Liu et al., 2024a), Logic-RL (Xie et al., 2025), cDPO (Lin et al., 2024) and GRPO (Shao et al., 2024). We select three models (Llama-3.2-3B (Grattafiori et al., 2024), Llama-3.1-8B (Grattafiori et al., 2024), Mistral-7B-v0.3 (Jiang et al., 2023)) as baselines.
@@ 5.1 Experimental Settings
5.2 Performance Comparison
(a) TruthfulQA Performance vs. Training Steps
(b) IFEval Performance vs. Training Steps
@@ 5.4 Additional Experiments
Case Study. A case study in Figure 3 on a medical query (see Appendix C.1 for details) demonstrates that, given the user prompt “I have a headache, what should I do?”, TI-DPO effectively assigns higher importance to safety-critical tokens (e.g., “medical attention”, “promptly”) in preferred responses, while penalizing risky suggestions (e.g., “painkillers”, “casually”) in non-preferred ones. Additionally, there are two other cases in Appendix C.2.
Figure 3: Case demo of responses to prompt “I have a headache, what should I do?”. Left: Preferred case. Middle: Intermediate case. Right: Non-preferred case. The darker color indicates higher weight.
Distribution of Weights. Figure 4 illustrates how TI-DPO dynamically adapts token importance weights based on task characteristics. For tasks relying on a few critical symbols (GSM8K, GPQA), weights concentrate in the [0.2, 0.5] range. Conversely, in tasks demanding strict adherence to safety or instructions (TruthfulQA, IFEval), weights shift to a higher [0.6, 0.8] interval. Meanwhile, for comprehensive tasks covering broader content (MMLU, HumanEval), the weight distribution is naturally more dispersed.
Figure 4: Distribution patterns of gradient-based token importance weights in six benchmark tasks (GSM8K, TruthfulQA, MMLU, GPQA, HumanEval, IFEval
Pearson correlation coefficient. To investigate the effectiveness of our hybrid weighting mechanism, we also conducted Pearson correlation coefficient analysis, with full results and methodology presented in Appendix B.1. Our analysis reveals a moderately strong positive correlation ($r\approx 0.65$) at the task level between average token importance and performance improvements, indicating that TI-DPO is particularly effective on tasks with concentrated weight distributions.
Robustness and Generation Diversity. We validate the robustness and generative diversity of TI-DPO, which can be seen in Appendix B.3. To evaluate the model’s stability, we tested accuracy under varying label noise levels. TI-DPO maintains superior performance compared to DPO and TPO as noise increases.
Sensitivity of Hyperparameters. We conducted sensitivity analyses for the weight-mixing parameter $\lambda$ and KL weight $\alpha$ in Appendix B.5. Performance remains stable for $\lambda\in[0.3,0.7]$. We have provided the specific values of the hyperparameters for this project, as shown in Appendix B.5.
@@ 6 Conclusion
We introduce TI-DPO, an optimization framework that effectively bridges the alignment gap between LLMs and human value systems. By introducing a mixed weight calculated collaboratively by gradient attribution and Gaussian prior, TI-DPO effectively overcomes the limitations of traditional DPO methods at the token level and their sensitivity to noise. On this basis, the triplet loss structure provides more refined guidance for model optimization. Theorem 2 and Theorem 3 theoretically illustrate the superiority of TI-DPO over DPO. The effectiveness of TI-DPO is unequivocally demonstrated through extensive experimentation. Our method achieves a state-of-the-art average score of 62.3 across a diverse suite of benchmarks, outperforming all baseline methods. As for the limitations, despite its effectiveness in fine-grained alignment, TI-DPO entails a computational overhead during training and performs slightly below sequence-level baselines on holistic reasoning tasks. Future work will focus on integrating our token-importance mechanism with group-based optimization methods like GRPO to bridge this gap and further enhance reasoning capabilities. More statements can be found in Appendix D.
Ethics Statement
This research focuses on aligning LLMs with human values to ensure beneficial, safe, and harmless interactions. We acknowledge that aligning models using human preference data carries the inherent risk of learning existing stereotypes or biases present in the training datasets. However, our proposed TI-DPO framework provides a structural advantage over standard methods by assigning explicit, interpretable importance weights to individual tokens. If the model reinforces a bias, these token weights make the source of the bias explicit and detectable, providing a direct mechanism for bias mitigation. All evaluations in this study were conducted on standard, publicly available benchmarks without involving human subject research.
Reproducibility Statement
To ensure full reproducibility of our work, the complete source code, training scripts, and demo are publicly available at https://github.com/gracefulning/TIDPO. We use exclusively open-weight base models (e.g., Llama-3.2-3B, Llama-3.1-8B, and Mistral-7B-v0.3) and public benchmark datasets. The paper provides comprehensive theoretical proofs and gradient derivations in Appendix A. Furthermore, the detailed implementation steps are clearly outlined in Algorithm 1, and all relevant hyperparameters used in our experiments are explicitly listed in Appendix B.5.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62301559. The authors would also like to thank the anonymous reviewers for their constructive feedback, as well as our colleagues for their insightful discussions during the development of the TI-DPO framework.
@@ A.4 Gradient Analysis of Loss
Appendix B Additional Experimental Results
We present some additional explanations and experimental results.
B.1 Pearson Correlation Coefficient
Table B1 presents multiple metrics for six tasks, namely GSM8K, GPQA, TruthfulQA, IFEval, MMLU, and HumanEval, including Q1, Q2 (median), Q3, average weight (mid-point of the interval), performance improvement $\Delta_{\text{ACC}}$ (%), and sample-level Pearson $r$ (Top 5 weights $vs$ accuracy). Among them, the “sample-level Pearson rr” is calculated based on the average weights of the Top-5 tokens for 100 questions in each task and the correctness of the answers to these questions (binary values), which reflects the microscopic internal correlation. The performance improvement $\Delta_{\text{ACC}}$ is calculated from the average accuracy gain of TI-DPO compared to DPO.
Table B1: Distribution of Token Importance Weight, Performance Improvement, and Sample-level Pearson Correlation Coefficient in Each Task
@@ B.1 Pearson Correlation Coefficient
The calculation process is divided into the calculation of microscopic and macroscopic Pearson correlation coefficients:
@@
B.2 Experimental Results Across Base Models
@@
B.3 Robustness and Generative Diversity
@@
B.4 Convergence Based on Theorem 2
@@
B.5 Sensitivity Analysis and Hyperparameters